Prove that n ! > 2^n for n a positive integer greater than or equal to 4.
Prove that LHS = RHS
1
Expert's answer
2020-05-13T19:06:53-0400

We apply the method of mathematical induction.

Induction basis : "n=4"



"\left\{\begin{array}{l}
n!=4!=1\cdot2\cdot3\cdot4=24\\
2^n=2^4=16
\end{array}\right.\longrightarrow 4!=24>16=2^4"

Induction assumption : suppose that the inequalities hold for all values "k\le n"



"k!>2^k, k=\overline{1,\ldots,n}"

Induction step : need to prove that



"\left(n+1\right)!>2^{n+1}"

Proof.



"\left(n+1\right)!=\left(n!\right)\cdot\left(n+1\right)>2^n\cdot\left(n+1\right)\to\\[0.3cm]
\left.2^n\cdot\left(n+1\right)>2^{n+1}\right|\div\left(2^n\right)\\[0.3cm]
n+1>2\to\\[0.3cm]
n>1-\text{true inequality, since by hypothesis}\,\,\,n>4"

Q.E.D.